# Distribution Backtracking Builds A Faster Convergence Trajectory for One-step Diffusion Distillation

### v1

@@ 4.3 Distribution Backtracking
$\text{Grad}(\eta)=E_{t,\epsilon}\left[s_{{\phi}}\left(\bm{x}_{t},t\right)-s_{\theta_{i}}^{\prime}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}$ (8)
@@ 5 Experiment
Experiments are conducted on different models across various datasets. We first compare the performance of DisBack with other multi-step diffusion models and distillation methods (Sec. 5.1). Secondly, we compare the convergence speed of DisBack with its variants without backtracking along the convergence trajectory (Sec. 5.2). Thirdly, further experiments are conducted to demonstrate DisBack’s effectiveness of mitigating the score mismatch issues (Sec. 5.3). Then, we also conduct the ablation study to show the effectiveness of DisBack (Sec. 5.4). Finally, we show the results of DisBack on text-to-image generation tasks (Sec. 5.5).
@@ 5.1 Quantitative evaluation
DisBack can achieve performance comparable to or even better than the existing diffusion models or distillation methods. Experiments are conducted on different diffusion models. (1) Both the teacher model and student generator are initialized by the pre-trained EDM model (Karras et al., 2022). (2) Both the teacher model and the student generator are initialized using a pre-trained CD model (Song et al., 2023). The performance of DisBack is shown in Tab. 1.
Table 1: The performance of DisBack on different datasets. The FID ($\downarrow$) scores are shown. DisBack + EDM/CD represents the DisBack method where both the teacher model and the student generator are initialized using a pre-trained EDM/CD.
Model NFE ($\downarrow$) FFHQ AFHQv2 ImageNet LSUN-bedroom LSUN-cat
DDPM (Ho et al., 2020) 1000 3.52 - 3.77 4.89 17.10
ADM (Dhariwal & Nichol, 2021) 1000 - - - 1.90 5.57
NSCN++ (Song et al., 2021b) 79 25.95 18.52 - - -
DDPM++ (Song et al., 2021b) 79 3.39 2.58 - - -
EDM (Karras et al., 2022) 79 2.39 1.96 2.44 3.57 6.69
EDM (Karras et al., 2022) 15 15.81 13.67 10.46 - -
PD (Salimans & Ho, 2021) 1 - - 8.95 16.92 29.60
CT (Song et al., 2023) 1 - - 13.00 16.00 20.70
CD (Song et al., 2023) 1 12.58 10.75 6.20 7.80 11.00
DisBack ($s_{\theta}$: EDM; $G_{stu}$: EDM) 1 10.88 - 8.44 - -
DisBack ($s_{\theta}$: CD; $G_{stu}$: CD) 1 11.99 9.97 5.73 6.99 10.30
Compared to multi-step generation diffusion models, the one-step EDM model distilled by the DisBack achieves comparable performance across different datasets. Specifically, it also outperforms the original EDM model with 15 NFEs (10.88 of DisBack and 15.81 of EDM on FFHQ64). Compared to existing one-step generators and distillation models, DisBack achieves optimal performance in most cases. Unlike consistency distillation, DisBack requires no training data and additional consistency constraints during training. Moreover, although CD already achieves excellent performance in one-step generation, distilling CD using DisBack can further enhance the performance of the original CD model, indicating that pre-trained models can benefit from DisBack. In conclusion, DisBack can achieve competitive distillation performance compared to existing models.
@@ 5.2 Convergence speed
We conducted a series of experiments on different models and datasets to demonstrate the advantages of DisBack in accelerating the convergence speed of the score distillation process: (1) The teacher model and the one-step student generator are both initialized by the pre-trained EDM (Karras et al., 2022) on unconditional FFHQ64 (Karras et al., 2019) and conditional ImageNet64 (Deng et al., 2009) datasets. (2) The teacher model and the one-step student generator are both initialized by the pre-trained DDIM (Song et al., 2021a) on unconditional CIFAR10 (Krizhevsky et al., 2009) datasets.
The results are shown in Fig. 1. When the teacher model and the one-step generator are both initialized by the EDM model, DisBack achieves a convergence speed 2.46 times faster than the variant without the backtracking strategy on the FFHQ64 dataset. On the ImageNet64 dataset, DisBack is 2.57 times faster. When both the teacher model and the one-step generator were initialized using the DDIM model, DisBack achieved a convergence speed 13.09 times faster than variants without backtracking. The fast convergence speed is because constraining the convergence trajectory of the generator provides a clear optimization direction, avoiding the generator falling into suboptimal solutions and enabling faster convergence to the target distribution.
@@ 5.3 Experiments on score mismatch issue
In this part, experiments are conducted to validate the positive impact of guiding the convergence trajectory and the effectiveness of mitigating the mismatch issues. We propose a new metric called mismatch degree to assess whether the predicted score of the pre-trained diffusion model matches the distribution’s real score given a data distribution. This score is inspired by the score-matching loss.
@@
Here $\bm{x}_{t}$ is the noisy data from the assessed distribution. Besides, $s_{\theta}(\bm{x}_{t},t)$ is the predicted score and $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the real score. We use Stable Target Field (STF) (Xu et al., 2022) to approximate the score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ of $x_{t}$ to the assessed distribution in experiments.
Figure 4: The mismatch degree during the distillation process of Diff-Instruct and proposed DisBack. The standard deviation is visualized. DisBack effectively mitigates the mismatch degree during the distillation process.
When the assessed distribution is close to the distribution represented by the teacher model $s_{\theta}$, the mismatch degree is small, and vice versa. When calculated directly on the training data, the resulting mismatch degree represents the ideal lower bound. Thus, the mismatch degree can be used to assess the convergence degree of the generated distribution and visualize the speed of convergence under the guidance of the convergence trajectory.
We conduct experiments on FFHQ dataset with Diff-Instruct (Luo et al., 2023b) as a baseline. We calculate the mismatch degree in the distribution of the student generator of both Diff-Instruct and DisBack. The pre-trained EDM model is chosen as the teacher model. In this scenario, the ideal lower bound is $0.037$. We visualized the mismatch degree in Fig. 4. With the guidance of the convergence trajectory from Disback, the mismatch degree is lower, meaning the student generator converges faster and better, achieving more effective distillation.
Table 2: Ablation study on introducing the convergence trajectory to the score distillation process. $s_{\theta}$ is the teacher model, $G_{stu}$ is the one-step generator. The FID ($\downarrow$) scores in each case are shown.
@@ 5.4 Ablation study
Ablation studies are conducted to compare the performance of DisBack with its variant without introducing the convergence trajectory. The results are shown in Tab. 2. Results show that the variant without the convergence trajectory constraint suffers from a performance decay in different cases. This confirms the efficacy of introducing the convergence trajectory between the teacher model and the student generator can improve the final performance of the generation.
@@
Model FID ($\downarrow$) NFE ($\downarrow$) User Preference
SDXL 19.36 100 38.7%
DisBack 18.96 1 61.3%
@@ 5.5 Text to Image Generation
Further experiments are conducted to introduce the convergence trajectory tasks. We use DisBack to distill the SDXL model (Podell et al., ) and evaluate the FID scores of the distilled SDXL and the original SDXL on the COCO 2014 (Radford et al., 2021). The user studies are conducted to further verify the effectiveness of DisBack. We randomly select 128 prompts from the LAION-Aesthetics (Schuhmann et al., 2022) to generate images and ask volunteer participants to choose the images they think are better. Detailed information about the user study is included in the Appendix B.3. The results of the FID evaluation and user study are presented in the table. DisBack achieved better results in single-step generation compared to the original SDXL with the 100-step DDIM sampler (Song et al., 2021a). The preference scores of DisBack over the original SDXL are $61.3\%$. Some generation samples are shown in Fig. 2 and Fig. 5.
We also conducted experiments on LCM-LoRA (Luo et al., 2023a) on the COCO 2014 (Radford et al., 2021) and show the efficacy of Disback. The LCM-LoRA distilled using DisBack has the FID score of 36.37, while the FID of LCM-LoRA is 78.26. Detailed experimental results are provided in the Appendix.
@@ 6 Conclusion
This paper proposes Distribution Backtracking Distillation (DisBack) to introduce the entire convergence trajectory of the teacher model in the score distillation. The DisBack can also be used in the distillation of large-scale text-to-image models. DisBack performs a faster and more efficient distillation and achieves a comparable or better performance in one-step generation compared to existing multi-step generation diffusion models and one-step diffusion distillation models.
@@
Although DisBack has produced excellent results in one-step generation, there is still some gap compared to advanced multi-step generation diffusion models. Moreover, the performance of DisBack is inherently limited by the teacher model. The better the original performance of the teacher model has, the better the performance of DisBack also presents. Additionally, to achieve optimal performances in both accelerated distillation and generation quality, DisBack requires careful design of the distribution degradation path and the setting of various hyperparameters (such as how many epochs used to fit each intermediate node). While with no meticulous design it can also achieve better performance than models without backtracking, further exploration is required to enable the model to reach optimal performance.
Appendix A More Experiment Results on Text-to-Image Generation
@@
Figure 6: The FID scores of LCM-LoRA distilled from SD1.5 across training steps. DisBack achieves faster convergence and better performance.
@@
We conducted the experiment on LCM-LoRA (Luo et al., 2023a). LCM-LoRA is a Low-Rank Adaptation (LoRA) fine-tuned version of the Latent Consistency Model (LCM), applicable across various fine-tuned Stable Diffusion models for high-quality, single-step or few-step generation. In this experiment, we choose LCM-LoRA as the student model and Stable Diffusion v1.5 as the teacher model. We observed that the score distillation underperforms when LCM-LoRA serves as the teacher model. This issue likely stems from the infeasibility of directly converting the outputs of LCM-LoRA into scores. We distill the LCM-LoRA on the LAION-Aesthetic-6.5+ dataset (Schuhmann et al., 2022) and evaluate the FID scores on the COCO 2014 dataset (Radford et al., 2021) at a resolution of 512×512. 50,000 real images and 30,000 generated images were used to calculate FID scores. The 30,000 generated images were obtained by generating one image for each of the 30,000 distinct prompts. The results are shown in Tab. 4. The change in FID scores over training steps is illustrated in Fig. 6, showing that our DisBack method achieves a 1.5 times acceleration in convergence speed and yields superior generation performance within the same training period. Some generated samples of DisBack and original LCM-LoRA are shown in Fig. 7.
Table 4: The distillation results of LCM-LoRA with NFE $=1$.
Model FID ($\downarrow$)
LCM-LoRA 78.26
DisBack 36.37

### v3

@@ 4.3 Distribution Backtracking
$\text{Grad}(\eta)=\mathbb{E}_{t,\bm{\epsilon}}\left[\left[s_{{\phi}}\left(\bm{x}_{t},t\right)-s_{\theta_{i}}^{\prime}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}\right]$ (8)
@@
Table 1: The unconditional generation performance of DisBack. The FID ($\downarrow$) scores are shown.
Model NFE ($\downarrow$) FFHQ AFHQv2 LSUN-bedroom LSUN-cat
DDPM (Ho et al., 2020) 1000 3.52 - 4.89 17.10
ADM (Dhariwal & Nichol, 2021) 1000 - - 1.90 5.57
NSCN++ (Song et al., 2021b) 79 25.95 18.52 - -
DDPM++ (Song et al., 2021b) 79 3.39 2.58 - -
EDM (Karras et al., 2022) 79 2.39 1.96 3.57 6.69
EDM (Karras et al., 2022) 15 15.81 13.67 - -
Diff-Instruct (Luo et al., 2023c) 1 19.93 - - -
PD (Salimans & Ho, 2021) 1 - - 16.92 29.60
CT (Song et al., 2023) 1 - - 16.00 20.70
CD (Song et al., 2023) 1 12.58 10.75 7.80 11.00
DisBack 1 10.88 9.97 6.99 10.30
@@ 5 Experiment
Experiments are conducted on different models across various datasets. We first compare the performance of DisBack with other multi-step diffusion models and distillation methods (Sec. 5.1). Secondly, we compare the convergence speed of DisBack with its variants without the constraint of the convergence trajectory (Sec. 5.2). Thirdly, further experiments are conducted to demonstrate DisBack’s effectiveness in mitigating the score mismatch issues (Sec. 5.3). Then, we also conduct the ablation study to show the effectiveness of introducing the convergence trajectory (Sec. 5.4). Finally, we show the results of DisBack on text-to-image generation tasks (Sec. 5.5).
Table 2: The conditional generation performance of DisBack on ImageNet 64x64 dataset.
Model NFE ($\downarrow$) FID ($\downarrow$)
DDPM (Ho et al., 2020) 1000 3.77
DDDM (Zhang et al., 2024) 1000 2.11
EDM (Karras et al., 2022) 511 1.36
EDM (Karras et al., 2022) 15 10.46
Moment Matching (Salimans et al., 2024) 8 3.3
SlimFlow (Zhu et al., 2024) 1 12.34
BOOT (Gu et al., 2024) 1 12.30
DDDM (Zhang et al., 2024) 1 3.47
CTM (Kim et al., 2024a) 1 2.06
Sid (Zhou et al., 2024) 1 1.52
DMD2 (Yin et al., 2024a) 1 1.51
Diff-Instruct (Luo et al., 2023c) 1 5.57
PD (Salimans & Ho, 2021) 1 8.95
CT (Song et al., 2023) 1 13.00
CD (Song et al., 2023) 1 6.20
DisBack 1 1.38
@@ 5.1 Quantitative evaluation
DisBack can achieve performance comparable to or even better than the existing diffusion models or distillation methods. Experiments are conducted on different datasets. (1) The unconditional generation on FFHQ 64x64, AFHQv2 64x64, LSUN-bedroom 256x256 and LSUN cat 256x256. (2) The conditional generation on ImageNet 64x64. The performance of DisBack is shown in Tab. 1 and Tab. 2. All the DisBack models are distilled from the pre-trained EDM model (Karras et al., 2022).
For unconditional generation, the one-step generator distilled by the DisBack achieves comparable performance across different datasets compared to multi-step generation diffusion models. Specifically, it outperforms the original EDM model with 15 NFEs (10.88 of DisBack and 15.81 of EDM on FFHQ64). Compared to existing one-step generators and distillation methods, DisBack achieves optimal performance. For conditional generation, the DisBack achieves the best performance compared to the existing models. Moreover, DisBack requires no training data and additional constraints during training. In conclusion, DisBack can achieve competitive distillation performance compared to existing models.
@@ 5.2 Convergence speed
We conducted a series of experiments to demonstrate the advantages of DisBack in accelerating the convergence speed of the score distillation process on unconditional CIFAR10 (Krizhevsky, 2009), FFHQ 64x64 (Karras et al., 2019), and conditional ImageNet 64x64 (Deng et al., 2009) datasets. Diff-Instruct (Luo et al., 2023c) is the existing SOTA score distillation method, which can be regarded as a variation of DisBack not introducing the convergence trajectory. We compared the FID trends of DisBack and Diff-Instruct during the distillation process in the same situation.
The results are shown in Fig. 1. As for unconditional generation, DisBack achieves a convergence speed 1.97 times faster than the variant without the constraint of convergence trajectory on the FFHQ 64x64 dataset and 5.67 times faster on the CIFAR10 dataset. For the conditional generation on the ImageNet 64x64 dataset, DisBack is 1.79 times faster than the variant without the constraint of convergence trajectory. The fast convergence speed is because constraining the convergence trajectory of the generator provides a clear optimization direction, avoiding the generator falling into suboptimal solutions and enabling faster convergence to the target distribution.
@@ 5.3 Experiments on score mismatch issue
In this part, experiments are conducted to validate the positive impact of constraining the convergence trajectory on mitigating the mismatch issues. We propose a new metric called mismatch degree to assess whether the predicted score of the teacher model matches the distribution’s real score given a data distribution. This score is inspired by the score-matching loss.
@@
Figure 4: The mismatch degree during the distillation process of Diff-Instruct and proposed DisBack. The standard deviation is visualized. DisBack effectively mitigates the mismatch degree during the entire distillation process.
Here $\bm{x}_{t}$ is the noisy data from the assessed distribution. Besides, $s_{\theta}(\bm{x}_{t},t)$ represents the predicted score and $\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ represents the real score. As shown in Eq. (4) and (6), the closer $s_{\theta}(\bm{x}_{t},t)$ is to $\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$, the more accurate the gradient estimation for the student generator obtains. Because the real scores are not available in practice, we use Stable Target Field (STF) (Xu et al., 2022) to approximate the real score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ on the assessed distribution. STF estimation leverages reference batches to reduce the variance of training objectives, which has been proven to yield accurate asymptotically unbiased estimates of the real score (Xu et al., 2022).
When the assessed distribution is close to the distribution of the teacher model $s_{\theta}$, the mismatch degree is small, and vice versa. When calculated directly on the training data, the resulting mismatch degree represents the ideal lower bound. Therefore, the mismatch degree can be used to assess the convergence degree of the generated distribution during the distillation process and visualize the convergence speed under the constraint of the convergence trajectory.
We conduct experiments on the FFHQ 64x64 dataset with Diff-Instruct (Luo et al., 2023c) as a baseline. We calculate the mismatch degree on the distribution of the student generator of both Diff-Instruct and the proposed DisBack. The pre-trained EDM model is chosen as the teacher model. In this scenario, the ideal lower bound of the mismatch degree is $0.037$. We visualized the mismatch degree in Fig. 4. With the constraining of the convergence trajectory, the mismatch degree of the proposed DisBack is lower during the distillation process, meaning the student generator converges faster and better. Thus, by constraining the convergence trajectory, the mismatch issue can be mitigated and DisBack can achieve more efficient distillation.
Table 3: Ablation study on constraining the convergence trajectory to the score distillation process. The FID ($\downarrow$) scores in each case are shown.
@@ 5.4 Ablation study
Ablation studies are conducted to compare the performance of DisBack with its variant without the constraint of the convergence trajectory. The results are shown in Tab. 3. Results show that the variant without the constraint of convergence trajectory suffers from a performance decay in different cases. This confirms the efficacy of constraining the convergence trajectory between the student generator and the teacher model can improve the final performance of the generation.
@@
Model NFE ($\downarrow$) FID ($\downarrow$) Patch-FID ($\downarrow$) CLIP ($\uparrow$)
LCM-SDXL Luo et al. (2023b) 1 81.62 154.40 0.275
LCM-SDXL Luo et al. (2023b) 4 22.16 33.92 0.317
DMD2 Yin et al. (2024a) 1 19.01 26.98 0.336
DMD2 Yin et al. (2024a) 4 19.32 20.86 0.332
SDXL-Turbo Sauer et al. (2025) 1 24.57 23.94 0.337
SDXL-Turbo Sauer et al. (2025) 4 23.19 23.27 0.334
SDXL Lightning Lin et al. (2024) 1 23.92 31.65 0.316
SDXL Lightning Lin et al. (2024) 4 24.46 24.56 0.323
SDXL Podell et al. (2024) 100 19.36 21.38 0.332
DisBack 1 18.96 26.89 0.335
@@ 5.5 Text to Image Generation
Further experiments are conducted on text-to-image generation tasks. We use DisBack to distill the SDXL model (Podell et al., 2024) and evaluate the FID scores of the distilled SDXL and the original SDXL on the COCO 2014 (Radford et al., 2021). The user studies are conducted to verify the effectiveness of DisBack. We randomly select 128 prompts from the LAION-Aesthetics (Schuhmann et al., 2022) to generate images and ask volunteer participants to choose the images they think are better. Detailed information about the user study is included in Sec. B.3. The results of the FID evaluation and user study are presented in Tab. 4. DisBack achieved optimal FID and comparable CLIP scores in single-step generation compared to baselines, but the Patch-FID showed a slight decay. The preference scores of DisBack over the original SDXL are $61.3\%$. Some generation samples are shown in Fig. 2 and Fig. 5.
We also conducted experiments on LCM-LoRA (Luo et al., 2023b). The LCM-LoRA distilled from SDv1.5 using DisBack has an FID score of 36.37 on one-step generation, while the FID score of the original LCM-LoRA is 78.26. Some generated samples of DisBack and original LCM-LoRA are shown in Fig. 6. The details of experiments and results are provided in Sec. A.1.
Figure 6: One step generation samples by original LCM-LoRA and its variant distilled from SD v1.5 with DisBack in 512×512. LCM-LoRA with DisBack can generate images with higher quality.
@@ 6 Conclusion
This paper proposes Distribution Backtracking Distillation (DisBack) to introduce the entire convergence trajectory of the teacher model in the score distillation. The DisBack can also be used to distill large-scale text-to-image models. DisBack performs a faster and more efficient distillation and achieves a comparable or better performance in one-step generation compared to existing multi-step generation diffusion models and one-step diffusion distillation models.
@@
The performance of DisBack is inherently limited by the teacher model. The better the original performance of the teacher model, the better the performance of DisBack will also be. Additionally, to achieve optimal performances in both accelerated distillation and generation quality, DisBack requires careful design of the distribution degradation path and the setting of various hyperparameters (such as how many epochs are used to fit each intermediate node in distribution backtracking stage). While with no meticulous design, it can also achieve better performance, further exploration is required to enable the model to reach optimal performance.
7 Acknowledgments
This work is supported by the Provincial Key Research and Development Plan of Zhejiang Province under No. 2024C01250(SD2), the National Key R&D Program of China under No. 2022YFB3303301, and the National Natural Science Foundation of China under No. 62006208. This work is supported by Alibaba-Zhejiang University Joint Research Institute of Frontier Technologies. The authors would like to thank anonymous reviewers for their helpful comments.
Appendix A More Experiment Results
A.1 Experiment Results on Text-to-Image Generation
@@
Figure 7: The FID scores of LCM-LoRA distilled from SD1.5 across training steps. The first 40 epochs refer to the computational overhead of the degradation recording stage of DisBack. DisBack achieves faster convergence and better performance.
@@
We conducted the experiment on LCM-LoRA (Luo et al., 2023b). LCM-LoRA is a Low-Rank Adaptation (LoRA) version of the Latent Consistency Model (LCM) Luo et al. (2023a), applicable across fine-tuned Stable Diffusion models for high-quality, single-step or few-step generation. In this experiment, we use LCM-LoRA as the student generator and Stable Diffusion v1.5 as the teacher model. We observed that the score distillation underperforms when LCM-LoRA serves as the teacher model. This issue likely stems from the infeasibility of directly converting the outputs of LCM-LoRA into scores.
We distill the LCM-LoRA with the proposed DisBack and evaluate the FID scores on the COCO 2014 dataset (Radford et al., 2021) with the resolution of 512×512. 50,000 real images and 30,000 generated images were used to calculate FID scores. The 30,000 generated images were obtained by generating one image for each of the 30,000 distinct prompts. In the case of one-step generation, the original LCM-LoRA has an FID score of $78.26$, while the DisBack achieves an FID of $36.37$. The change in FID scores over training steps is illustrated in Fig. 7, showing that DisBack achieves a 1.5 times acceleration in convergence speed and yields superior generation performance within the same training period.
A.2 Additional Experiment Results
To explore the distillation performance of the proposed method when the architectures of the teacher and student models differ, we opt for EDM Karras et al. (2022) as the pre-trained diffusion model, and FastGAN Liu et al. (2020) architecture for the student model to conduct the experiment. Table 5 shows the performance of baselines and the proposed models on the distillation task from a diffusion model to a generator. Compared to the original FastGAN, DisBack can effectively improve the generation quality. The results also show that the one-step sampling performance of DisBack is better than Score GAN and EDM with 11 NFEs.
Table 5: The performance of DisBack on the distillation from pre-trained EDM model to FastGAN on FFHQ, AFHQv2, and CelebA in the resolution of $64\times 64$.
Model FFHQ AFHQv2 CelebA
NFE ($\downarrow$) FID ($\downarrow$) IS (↑) FID ($\downarrow$) IS (↑) FID ($\downarrow$) IS (↑)
FastGAN (Liu et al., 2020) 1 30.27 2.37 28.59 5.94 29.35 2.36
EDM (NFE 11) (Karras et al., 2022) 11 29.28 2.97 13.67 10.86 23.05 3.01
Score GAN (Franceschi et al., 2023) 1 43.89 2.13 53.86 2.13 50.41 2.12
DisBack 1 23.84 3.27 18.95 7.00 23.16 3.02
A.3 Visualization of Intermediate Teacher Trajectory
To further demonstrate the effectiveness of the degradation path, we visualized the images generated by the initial generator, the intermediate checkpoints and the teacher model, along the degradation path. As shown in Fig. 8 and 9, We can observe that the images generated by the first node in the trajectory are similar to those of the initial generator, while the images generated by the last node in the trajectory are close to those of the teacher model. This is consistent with our theoretical analysis.